OpenStudy (anonymous):

Mike shoots a large marble (Marble A, mass: 0.05 kg) at a smaller marble (Marble B, mass: 0.03 kg) that is sitting still. Marble A was initially moving at a velocity of 0.6 m/s, but after the collision it has a velocity of –0.2 m/s. What is the resulting velocity of marble B after the collision?

OpenStudy (anonymous):

use these equations to solve \[v1=\frac{ (m1-m2)u1 }{ m1+m2 }+\frac{ 2*m2u2 }{ m1+m2 }\] and \[v2=\frac{ 2m1u1 }{ m1+m2 }+\frac{(m2-m1)u2 }{m1+m2 }\] wher m1 = mass of body A m2=mass of body B u1 =initial velocity of body A u2=initial velocity of bodyB v1=final velocity that is velocity after collision for body A v2=final velocity that is velocity after collision for body B